OpenStudy (anonymous):

a planet of mass M has a satellite of mass m, located a distance d away, attracted by a force F. if the mass of the satellite is double while d and M remain unchanged by what factor will F change?

OpenStudy (anonymous):

take a simple way General formula of Newton's Graviity is F = (G M m)/ d^2 where: F ~ M F ~ m F ~ 1/d^2 G is gravity constanta so unchange, and if the mass of the satellite is double while d and M remain unchanged So, the factor will change of F is F ~ M with M1 = M, M2 = 2M, F1 = F, and F2 =? \[\frac{ F _{1} }{ F _{2}} = \frac{ M _{1} }{M _{2} }\] So \[F _{2} = F _{1} \times \frac{ M _{2} }{ M _{1} } = F \times \frac{ 2M }{ M } = 2F\]

OpenStudy (anonymous):

SO if the mass of the satellite is double the force of gravity is double!!!

OpenStudy (anonymous):

\[F = \frac{ GMm }{ r ^{2} }\] So yes, if you double the mass (small m) and leave everything else the same, then you double F, the force due to gravity.

OpenStudy (anonymous):

nice @furnessj :)